OpenStudy (anonymous):

Write the standard form of the equation of line passing through the point (-1,-3) and perpendicular to the line x - 4y = 12 1) -4 -y = -7 2) 4x+y = -7 3) x + 4y = -13 4) none of the above

OpenStudy (anonymous):

x-4y = 12 y = (x/4) - 3 slope of this line is 1/4 but we want line perpendicular to this one so you take the negative reciprocal : -4 now line equation : y = y1 + m(x-x1) y1 = -3 x1 = -1 m = -4 y = -3 + -4(x+1) y = -3 -4x -4 y = -4x -7 which is 2 since 4x +y = -7 -> y = -4x -7